# Decision Mamba: A Multi-Grained State Space Model with Self-Evolution Regularization for Offline RL

### v2

@@ Abstract
While the conditional sequence modeling with the transformer architecture has demonstrated its effectiveness in dealing with offline reinforcement learning (RL) tasks, it is struggle to handle out-of-distribution states and actions. Existing work attempts to address this issue by data augmentation with the learned policy or adding extra constraints with the value-based RL algorithm. However, these studies still fail to overcome the following challenges: (1) insufficiently utilizing the historical temporal information among inter-steps, (2) overlooking the local intra-step relationships among states, actions and return-to-gos (RTGs), (3) overfitting suboptimal trajectories with noisy labels. To address these challenges, we propose Decision Mamba (DM), a novel multi-grained state space model (SSM) with a self-evolving policy learning strategy. DM explicitly models the historical hidden state to extract the temporal information by using the mamba architecture. To capture the relationship among state-action-RTG triplets, a fine-grained SSM module is designed and integrated into the original coarse-grained SSM in mamba, resulting in a novel mamba architecture tailored for offline RL. Finally, to mitigate the overfitting issue on noisy trajectories, a self-evolving policy is proposed by using progressive regularization. The policy evolves by using its own past knowledge to refine the suboptimal actions, thus enhancing its robustness on noisy demonstrations. Extensive experiments on various tasks show that DM outperforms other baselines substantially. Our code is available at  https://github.com/aopolin-lv/DecisionMamba.
@@ 1 Introduction
Offline Reinforcement Learning (RL) [13, 27, 29, 38] has attracted great attention due to its remarkable successes in the fields of robotic control [5, 36] and games [3, 32, 50]. As transformer [49] has exhibited powerful sequential modeling abilities in natural language processing [4, 43] and computer vision [10, 42], many efforts [6, 8, 25, 61] have been made on applying this architecture to offline RL tasks. Transformer-based methods view the state, action, and reward/return-to-go (RTG) as a sequence, and then predict actions by using the transformer encoder. However, it often fails to make correct decisions when encountering out-of-distribution states or actions, showing limited robustness. Previous work attempts to address this issue from the perspective of data augmentation [51, 64] and objective constraints [6, 53, 61]. However, they introduce a significant number of noises or the overestimation bias. Thus, how to enhance model robustness remains a highly challenging and insufficiently explored issue.
@@
In this study, we offer two novel perspectives on improving model robustness through both the model architecture and learning strategy. In terms of the model architecture, (1) although previous studies have made some modifications to the transformer architecture [23, 44, 52], they have not fully utilized inter-step information, particularly historical information which is critical for decision-making processes. For example, the robot can adjust its subsequent routes based on the historical information of failed paths for completing the navigation task; (2) furthermore, most existing approaches adopt transformer to model the flattened trajectory as a sequence, while ignoring the structural trajectory patterns of the causal intra-step relationship among states, actions, and RTGs (SARs). A RL policy typically predicts the next action given the current state based on the RTG. Thus, this kind of fine-grained intrinsic connection among SARs is intuitively beneficial for policy learning. As regards to the learning strategy, (3) there exists a large number of noisy labels in the suboptimal trajectories which hurt the performance of the policy significantly. Although the existing work that generates pseudo trajectories or actions alleviates this problem to some extent [57, 64], it also introduces other biases or errors.
@@
To address the above issues, we propose Decision Mamba (DM), a multi-grained state space model with a self-evolving policy learning strategy for offline RL. In order to adequately leverage the historical information, we adopt mamba to explicitly model the temporal state among inter-steps, since mamba architecture [15, 18, 40] shows a more effective capability of extracting the historical information. Meanwhile, the causal intra-step relationship is beneficial for the model to understand the common patterns within the local dynamics. Thus, we introduce a fine-grained SSM module to extract the local features of structural patterns among the SAR triplet within each intra-step. Apart from modifying the model architecture and aligning it to the trajectory pattern, we also propose a learning strategy to prevent the policy from overfitting noisy labels. This is achieved by a progressive self-evolution regularization which leverages the past knowledge of the policy itself to refine and adjust the target label adaptively.
@@ 3.2.1 Multi-Grained Mamba
Following the sequence modeling, we first use multilayer perceptrons (MLPs) to embed the SARs from the given trajectory $\tau=(s_{0},R_{0},a_{0},\ldots,s_{T},R_{T},a_{T})$. Then, the trajectory embeddings are added the absolute step position embeddings to attach the position information, similar to the classic usage in the NLP field. Mathematically, it can be formulated as follows:
$e^{\mathrm{s}}_{i}=\operatorname{MLP}(s_{i}),\ \ \ \ \ \ e^{\mathrm{R}}_{i}=\operatorname{MLP}(R_{i}),\ \ \ \ \ \ e^{\mathrm{a}}_{i}=\operatorname{MLP}(a_{i}),$ (8)
$e_{i}=[e^{\mathrm{s}}_{i};e^{\mathrm{R}}_{i};e^{\mathrm{a}}_{i}]+\operatorname{broadcast}(e^{\mathrm{t}}_{i}),$ (9)
@@
0: trajectory sequence $\tau=(s_{0},R_{0},a_{0},\ldots,s_{t},R_{t})$
@@
2:$\mathbf{s},\mathbf{R},\mathbf{a}$: $(\mathtt{B},\mathtt{L})$leftarrow $\mathrm{Split}(\tau_{t-l:t})$
3:$\mathbf{e}^{\mathbf{s}},\mathbf{e}^{\mathbf{R}},\mathbf{e}^{\mathbf{a}}$: $(\mathtt{B},\mathtt{L},\mathtt{D})$leftarrow $\mathrm{MLP}(\textbf{s}),\mathrm{MLP}(\textbf{R}),\mathrm{MLP}(\textbf{a})$
4:$\mathbf{h}_{0}$: $(\mathtt{B},\mathtt{L},\mathtt{D})$leftarrow $\mathrm{Flatten}(\mathbf{e}^{\mathbf{s}},\mathbf{e}^{\mathbf{R}},\mathbf{e}^{\mathbf{a}})$
@@
Further, to better discern the dependencies among SARs within each step, we gather the feature of each single step to obtain the fine-grained representation via a 1D-convolution layer, and then introduce a fine-grained SSM module for extracting the local pattern among SAR, as shown in the middle part and right part of Figure 1.
@@ 3.2.2 Progressive Self-Evolution Regularization
$\mathop{\mathrm{minimize}}\limits_{\theta}\ \mathbb{E}_{s_{t},\tau\sim D_{\mu}}\big{[}\log\pi_{\theta}(\tilde{a}_{t}|s_{t},R_{t},\tau_{<t})\big{]}.$ (21)
@@ 3.2.3 Training Objective
To make the training procedure more robust, we introduce the inverse training goals: predicting the next state and the next RTG. Individuals often assess the feasibility of actions by envisioning their potential outcomes. Therefore, we expect the policy to predict the post-execution state and RTG based on the predicted action, thus improving its robustness. Specifically, given the trajectory $\tau=(s_{0},R_{0},a_{0},\ldots,s_{t},R_{t})$, Decision Mamba originally predicts the next action $\hat{a}_{t}$. Further, by incorporating the action $a_{t}$ to the original trajectory $\tau_{t-l:t}$, it is also predicts the next state $\hat{s}_{t}$ and the next RTG $\hat{R}_{t}$. Compared to the Eq. (3), the training objective of DM with the refined target can be written as follows:
$\mathop{\mathrm{minimize}}\limits_{\theta}\ \mathbb{E}_{s_{t},\tau\sim D_{\mu}}\Big{[}\sum_{t=0}^{T}\big{[}\lambda_{1}\underbrace{\log\pi_{\theta}(\tilde{a}_{t}|s_{t},R_{t},\tau_{<t})}_{\mathrm{PSER}}+\lambda_{2}\underbrace{\log\pi_{\theta}(s_{t}|\tau_{<t})}_{\mathrm{predicting\ states}}+\lambda_{3}\underbrace{\log\pi_{\theta}(R_{t}|\tau_{<t})}_{\mathrm{predicting\ RTGs}}\big{]}\Big{]},$ (23)
@@ A.2 Implement Details
Antmaze 1× $1\ \times\ 10^{-4}$ 1× $1\ \times\ 10^{-5}$ $20$ $5,20$ 100K $0.95$ $0.85$

### v3

@@ Abstract
While the conditional sequence modeling with the transformer architecture has demonstrated its effectiveness in dealing with offline reinforcement learning (RL) tasks, it is struggle to handle out-of-distribution states and actions. Existing work attempts to address this issue by data augmentation with the learned policy or adding extra constraints with the value-based RL algorithm. However, these studies still fail to overcome the following challenges: (1) insufficiently utilizing the historical temporal information among inter-steps, (2) overlooking the local intra-step relationships among return-to-gos (RTGs), states and actions, (3) overfitting suboptimal trajectories with noisy labels. To address these challenges, we propose Decision Mamba (DM), a novel multi-grained state space model (SSM) with a self-evolving policy learning strategy. DM explicitly models the historical hidden state to extract the temporal information by using the mamba architecture. To capture the relationship among RTG-state-action triplets, a fine-grained SSM module is designed and integrated into the original coarse-grained SSM in mamba, resulting in a novel mamba architecture tailored for offline RL. Finally, to mitigate the overfitting issue on noisy trajectories, a self-evolving policy is proposed by using progressive regularization. The policy evolves by using its own past knowledge to refine the suboptimal actions, thus enhancing its robustness on noisy demonstrations. Extensive experiments on various tasks show that DM outperforms other baselines substantially.
footnotetext: Corresponding Author.
@@ 1 Introduction
Offline Reinforcement Learning (RL) [13, 27, 29, 38] has attracted great attention due to its remarkable successes in the fields of robotic control [5, 36] and games [3, 32, 50]. As transformer [49] has exhibited powerful sequential modeling abilities in natural language processing [4, 43] and computer vision [10, 42], many efforts [6, 8, 25, 61] have been made on applying this architecture to offline RL tasks. Transformer-based methods view the reward/return-to-go (RTG), state, and action as a sequence, and then predict actions by using the transformer encoder. However, it often fails to make correct decisions when encountering out-of-distribution states or actions, showing limited robustness. Previous work attempts to address this issue from the perspective of data augmentation [51, 64] and objective constraints [6, 53, 61]. However, they introduce a significant number of noises or the overestimation bias. Thus, how to enhance model robustness remains a highly challenging and insufficiently explored issue.
@@
In this study, we offer two novel perspectives on improving model robustness through both the model architecture and learning strategy. In terms of the model architecture, (1) although previous studies have made some modifications to the transformer architecture [23, 44, 52], they have not fully utilized inter-step information, particularly historical information which is critical for decision-making processes. For example, the robot can adjust its subsequent routes based on the historical information of failed paths for completing the navigation task; (2) furthermore, most existing approaches adopt transformer to model the flattened trajectory as a sequence, while ignoring the structural trajectory patterns of the causal intra-step relationship among RTGs, states, and actions (RSAs). A RL policy typically predicts the next action given the current state based on the RTG. Thus, this kind of fine-grained intrinsic connection among RSAs is intuitively beneficial for policy learning. As regards to the learning strategy, (3) there exists a large number of noisy labels in the suboptimal trajectories which hurt the performance of the policy significantly. Although the existing work that generates pseudo trajectories or actions alleviates this problem to some extent [57, 64], it also introduces other biases or errors.
@@
To address the above issues, we propose Decision Mamba (DM), a multi-grained state space model with a self-evolving policy learning strategy for offline RL. In order to adequately leverage the historical information, we adopt mamba to explicitly model the temporal state among inter-steps, since mamba architecture [15, 18, 40] shows a more effective capability of extracting the historical information. Meanwhile, the causal intra-step relationship is beneficial for the model to understand the common patterns within the local dynamics. Thus, we introduce a fine-grained SSM module to extract the local features of structural patterns among the RSA triplet within each intra-step. Apart from modifying the model architecture and aligning it to the trajectory pattern, we also propose a learning strategy to prevent the policy from overfitting noisy labels. This is achieved by a progressive self-evolution regularization which leverages the past knowledge of the policy itself to refine and adjust the target label adaptively.
@@ 3.2.1 Multi-Grained Mamba
Following the sequence modeling [8], we first use multilayer perceptrons (MLPs) to embed the RSAs from the given trajectory $\tau=(R_{0},s_{0},a_{0},\ldots,R_{T},s_{T},a_{T})$. Then, the trajectory embeddings are added the absolute step position embeddings to attach the position information, similar to the classic usage in the NLP field. Mathematically, it can be formulated as follows:
$e^{\mathrm{R}}_{i}=\operatorname{MLP}(R_{i}),\ \ \ \ \ \ e^{\mathrm{s}}_{i}=\operatorname{MLP}(s_{i}),\ \ \ \ \ \ e^{\mathrm{a}}_{i}=\operatorname{MLP}(a_{i}),$ (8)
$e_{i}=[e^{\mathrm{R}}_{i};e^{\mathrm{s}}_{i};e^{\mathrm{a}}_{i}]+\operatorname{broadcast}(e^{\mathrm{t}}_{i}),$ (9)
@@
0: trajectory sequence $\tau=(R_{0},s_{0},a_{0},\ldots,R_{t},s_{t})$
@@
2:$\mathbf{R},\mathbf{s},\mathbf{a}$: $(\mathtt{B},\mathtt{L})$leftarrow $\mathrm{Split}(\tau_{t-l:t})$
3:$\mathbf{e}^{\mathbf{R}},\mathbf{e}^{\mathbf{s}},\mathbf{e}^{\mathbf{a}}$: $(\mathtt{B},\mathtt{L},\mathtt{D})$leftarrow $\mathrm{MLP}(\textbf{R}),\mathrm{MLP}(\textbf{s}),\mathrm{MLP}(\textbf{a})$
4:$\mathbf{h}_{0}$: $(\mathtt{B},\mathtt{L},\mathtt{D})$leftarrow $\mathrm{Flatten}(\mathbf{e}^{\mathbf{R}},\mathbf{e}^{\mathbf{s}},\mathbf{e}^{\mathbf{a}})$
@@
Further, to better discern the dependencies among RSA within each step, we gather the feature of each single step to obtain the fine-grained representation via a 1D-convolution layer, and then introduce a fine-grained SSM module for extracting the local pattern among RSA, as shown in the middle part and right part of Figure 1.
@@ 3.2.2 Progressive Self-Evolution Regularization
$\mathop{\mathrm{minimize}}\limits_{\theta}\ \mathbb{E}_{s_{t},\tau\sim D_{\mu}}\big{[}\log\pi_{\theta}(\tilde{a}_{t}|R_{t},s_{t},\tau_{<t})\big{]}.$ (21)
@@ 3.2.3 Training Objective
To make the training procedure more robust, we introduce the inverse training goals: predicting the next state and the next RTG. Individuals often assess the feasibility of actions by envisioning their potential outcomes. Therefore, we expect the policy to predict the post-execution state and RTG based on the predicted action, thus improving its robustness. Specifically, given the trajectory $\tau=(R_{0},s_{0},a_{0},\ldots,R_{t},s_{t})$, Decision Mamba originally predicts the next action $\hat{a}_{t}$. Further, by incorporating the action $a_{t}$ to the original trajectory $\tau_{t-l:t}$, it is also predicts the next RTG $\hat{R}_{t+1}$ and the next state $\hat{s}_{t+1}$. Compared to the Eq. (3), the training objective of DM with the refined target can be written as follows:
$\mathop{\mathrm{minimize}}\limits_{\theta}\ \mathbb{E}_{s_{t},\tau\sim D_{\mu}}\Big{[}\sum_{t=0}^{T}\big{[}\lambda_{1}\underbrace{\log\pi_{\theta}(\tilde{a}_{t}|R_{t},s_{t},\tau_{<t})}_{\mathrm{PSER}}+\lambda_{2}\underbrace{\log\pi_{\theta}(R_{t+1}|\tau_{\leq t})}_{\mathrm{predicting\ RTGs}}+\lambda_{3}\underbrace{\log\pi_{\theta}(s_{t+1}|\tau_{\leq t})}_{\mathrm{predicting\ states}}\big{]}\Big{]},$ (23)
@@ A.2 Implement Details
Antmaze 1× $1\ \times\ 10^{-4}$ 1× $1\ \times\ 10^{-5}$ $20$ $5,20$ 100K $0.95$ $0.5$